Chronos Time Pieces. Chronos Time Pieces of Boston exports watches to many countries, selling in local currencies to stores and distributors. Chronos prides itself on being financially conservative. At least 70% of each individual transaction exposure is hedged, mostly in the forward market, but occasionally with options. Chronos' foreign exchange policy is such that the 70% hedge may be increased up to a 120% hedge if devaluation or depreciation appears imminent. Chronos has just shipped to its major North American distributor. It has issued a 90-day invoice to its buyer for euro€1 comma 570 comma 0001,570,000. The current spot rate is $1.22281.2228 /euro€, the 90-day forward rate is $1.2272 /euro€. Chronos' treasurer, Manny Hernandez, has a very good track record in predicting exchange rate movements. He currently believes the euro will weaken against the dollar in the coming 90 to 120 days, possibly to around $1.1602 /euro€.

a. Case 1: Manny is right and the spot rate in 90 days is $1.1602 /euro€.

1. How much in U.S. dollars will Chronos receive in 90 days if 100% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.)

2. How much in U.S. dollars will Chronos receive in 90 days if 70% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.)

3. How much in U.S. dollars will Chronos receive in 90 days if 120% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.)

Case 2: Manny is wrong and the spot rate in 90 days is $1.2665 /euro€.

1. How much in U.S. dollars will Chronos receive in 90 days if 100% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.) 2.How much in U.S. dollars will Chronos receive in 90 days if 70% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.) 3.How much in U.S. dollars will Chronos receive in 90 days if 120% of the transaction exposure is hedged with the forward contract? $ (Round to the nearest dollar.)
Answer #1
a) Case 1:
1) Amount to be received in 90 days = 1570000*1.2272 = $        19,26,704
2) Amount to be received in 90 days under the forward contract = 1570000*1.2272*70% = $        13,48,693
Balance to be received at spot rate = 1570000*30%*1.1602 = $          5,46,454
Total amount to be received in 90 days $        18,95,147
3) Amount to be received in 90 days against receivables = 1570000*1.2272*100% = $        19,26,704
Net amount realized from balance to be sold at the spot rate = 1570000*20%*(1.1602-1.2272) = $            -21,038
Total amount to be received in 90 days $        19,05,666
Case 2:
1) Amount to be received in 90 days = 1570000*1.2272 = $        19,26,704
2) Amount to be received in 90 days under the forward contract = 1570000*1.2272*70% = $        13,48,693
Balance to be received at spot rate = 1570000*30%*1.2665 = $          5,96,522
Total amount to be received in 90 days $        19,45,214
3) Amount to be received in 90 days against receivables = 1570000*1.2272*100% = $        19,26,704
Net amount realized from balance to be sold at the spot rate = 1570000*20%*(1.2665-1.2272) = $              12,340
Total amount to be received in 90 days $        19,39,044